OpenStudy (anonymous):

Can someone help me with implicit differentiation please: (xy)^1/2 + 2x = y^1/2

OpenStudy (michele_laino):

using the cahin rule, and the Leibniz rule, we can write this: \[\Large \frac{{d\sqrt {xy} }}{{dx}} = \left( {\sqrt {xy} } \right)' = \frac{1}{2}\frac{1}{{\sqrt {xy} }}\left( {y + xy'} \right)\]

OpenStudy (michele_laino):

oops..chain*

OpenStudy (michele_laino):

being y=y(x), namely y is a function of x

OpenStudy (anonymous):

I am a bit lost on how you get (y+xy')

OpenStudy (michele_laino):

since we hav to compute the first derivative of this function: \[\Large {xy}\]

OpenStudy (michele_laino):

using the product rule or Leibniz rule

OpenStudy (michele_laino):

have*

OpenStudy (anonymous):

ahhhhh I got it thank you :)

OpenStudy (michele_laino):

thank you! :)

OpenStudy (anonymous):

then is it + 2

OpenStudy (michele_laino):

if we differentiate all of your terms, we get: \[\Large \frac{1}{2}\frac{1}{{\sqrt {xy} }}\left( {y + xy'} \right) + 2 = \frac{1}{2}\frac{1}{{\sqrt y }}y'\]

OpenStudy (michele_laino):

now, the least common multiple is: \[\Large \sqrt {xy} = \sqrt x \sqrt y \]

OpenStudy (michele_laino):

so we have to multiply both sides of that expression, by : \[\Large\sqrt x \sqrt y \]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

1/2(y+xy')+2=x^2sqrty/2 y' ??

OpenStudy (michele_laino):

I got this: \[\Large \frac{1}{2}\left( {y + xy'} \right) + 2\sqrt {xy} = \frac{1}{2}\sqrt x y'\]

OpenStudy (anonymous):

yes ok I understand that :)

OpenStudy (michele_laino):

really, the least common multiple was \[\Large 2\sqrt {xy} = 2\sqrt x \sqrt y \] nevertheless never mind, since we can multiply, both sides of last expression, by 2 right now

OpenStudy (michele_laino):

so, if we multiply both sides of my last expression by 2, we get: \[\Large y + xy' + 4\sqrt {xy} = \sqrt x y'\]

OpenStudy (anonymous):

yes I got that as well :)

OpenStudy (michele_laino):

now we subtract, from both sides this quantity: \[\Large xy'\] so we can write: \[\Large y + 4\sqrt {xy} = \sqrt x y' - xy'\]

OpenStudy (anonymous):

then take out the y' as the common factor?

OpenStudy (michele_laino):

better is to factor out: \[\Large \sqrt x y'\] at the right side

OpenStudy (michele_laino):

so we can write: \[\Large y + 4\sqrt {xy} = \sqrt x y'\left( {1 - \sqrt x } \right)\]

OpenStudy (anonymous):

then divide the other side by sort x(1-sqrtx)?

OpenStudy (anonymous):

oops not sort sqrt

OpenStudy (michele_laino):

yes! we have to divide both sides, by: \[\Large \sqrt x \left( {1 - \sqrt x } \right)\]

OpenStudy (michele_laino):

so we can write the expression for y', what do you get?

OpenStudy (anonymous):

\[y'=\frac{ y+4\sqrt{xy} }{ (\sqrt{x} -x)}\]

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Are you able to walk me through a couple more if you have time, I am still struggling to get my head around them :)

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

Thank you so very much I really appreciate it, I can close this and put up a new one so I can give you another medal :)

OpenStudy (michele_laino):

ok! :)